OpenStudy (anonymous):

If f(x)=(x=8)^2, use the difference quotient to find Ax+Bh+C. A=2 B=1 C=?

OpenStudy (anonymous):

Hi, re-write the f(x), (x=8)^2 means?

OpenStudy (anonymous):

f(x)=(x+8)^2

OpenStudy (anonymous):

sorry about that

OpenStudy (anonymous):

expand followed the formula of (a +b)^2 = a^2 +2ab +b^2. your a is x, your b is 8. then you get .... Let me know the result. and then i guide you more

OpenStudy (anonymous):

\[\frac{ f(x+h)-f(x) }{ h }= Ax+ Bh+ C\] A=2 B=1

OpenStudy (anonymous):

that is the formula for limit, right?

OpenStudy (anonymous):

I have C=1/16. what do you think?

OpenStudy (anonymous):

oh, sorry, I make mistake, not that. a moment

OpenStudy (anonymous):

okay, i've been working on this forever. no it's difference quotient. 1/16 is incorrect. I got zero but that's wrong as well

OpenStudy (anonymous):

wait a moment , i nearly got it.

OpenStudy (anonymous):

please, let check whether C =14

OpenStudy (anonymous):

No, it must be C=16

OpenStudy (anonymous):

\[f(x) = (x+8)^2 = x^2 +16x+64 \[f(x+h) = (x+h+8)^2 = x^2 +2xh +16x+h^2+16h+64\]

OpenStudy (anonymous):

combine them into the your formula to get \[\frac{ 2xh+h^2+16h }{h}= 2x+h+16\]

OpenStudy (anonymous):

you have A=2; B=1 and C=16. Hope this is right

OpenStudy (anonymous):

that's correct. thanks a ton